OpenStudy (anonymous):

What does the Pythagorean Theorem state about the sides of a right triangle if the legs of triangle have the measurement of a and b and the hypotenuse of the triangle measures c?

OpenStudy (anonymous):

http://www.mathsisfun.com/pythagoras.html

OpenStudy (anonymous):

Possible answers are A. a2 − b2 = c2 B. a + b + c C. a2 + b2 = c2 D. a3 + b3 = c3

OpenStudy (anonymous):

Read Up!

OpenStudy (anonymous):

\[c^2 = a^2 + b^2\]